RAMAN STUDY OF MOLECULAR VIBRATIONAL RELAXATION IN H-BONDED SYSTEMS

\underline{V. POGORELOV}, E. HODJIEVA, A. LIZENGEVICH, A. VINIJCHUK, Department of Physics, Kiev Taras Shevchenko’s University, Glushkov pr. 6, Kiev 252127, Ukraine.

The idea of the paper is a comparative study of intermolecular H-bond interaction influence on a shape and a shift of bands wich correspond to vibrations of two types: 1)vibrations wich are subjected to H-bond influence directly; 2)vibrations wich are affected by H-bond indirectly. A study of feature of H-bond influence on a shape of Fermi-duplicates was perfopmed too. Solvents with weak, intermediate and strong H-bond were selected. Principal results: 1)There isn’t direct correlation between energy of specific interaction and shifts of CH-bands. 2)Intramolecular vibrational redistribution is weakly subjected to influence of weak and intermediate specific interactions. In the case of a strong H-bond intermolecular vibratinal exchange predominates over intramolecular process.